OpenStudy (anonymous):

Prove: Cot^2x-Csc^2x

OpenStudy (mertsj):

Prove what?

OpenStudy (anonymous):

thats all my teacher wrote.

OpenStudy (mertsj):

Did she write "prove" or "simplify"

OpenStudy (anonymous):

Probably Simplify

OpenStudy (mertsj):

Then replace cot with cos/sin and replace csc with 1/sin

OpenStudy (anonymous):

ok, is that how far is simplifs

OpenStudy (mertsj):

Then add the fractions.

OpenStudy (anonymous):

now i have CosX-1/SinX

OpenStudy (mertsj):

\[\cot ^2-\csc ^2=\frac{\cos ^2}{\sin ^2}-\frac{1}{\sin ^2}\]

OpenStudy (anonymous):

Sin^2/Sin^2=1

OpenStudy (mertsj):

\[\frac{\cos ^2x-1}{\sin ^2x}=\frac{-1(1-\cos ^2x)}{\sin ^2x}=\frac{-1(\sin ^2x)}{ \sin ^2x}=-1\]

OpenStudy (anonymous):

ok, thank you.

OpenStudy (mertsj):

yw

OpenStudy (mertsj):

Or if you have had the Pythagorean identities, you could use those.

OpenStudy (anonymous):

yeah, proofs are just confusing.